Automorphisms of the Computably Enumerable Sets

Leo Harrington

There is a technique for building Δ_{3}^{0} automorphisms of the lattice E of computably enumerable sets. A recent result with Peter Cholak shows that there are certain parts of E (actually these are parts of E^{eq}) such that any automorphism restricted to such parts must be Δ_{3}^{0}. This talk will be about this result and some of its consequences.

Address: Department of Mathematics, University of California, Berkeley, CA 94720-3840, USA

E-mail: leo@math.berkeley.edu